OpenStudy (anonymous):

State the first 5 terms of the arithmetic sequence if t1=1 and d=4.

OpenStudy (butterflyhope):

Well, if a1 = 1 and the difference is 4, what's 1 + 4? That's your a2. a1 = 1. a2 = 5. a3 = so on and so forth...

OpenStudy (anonymous):

oh okay thanks

OpenStudy (butterflyhope):

Do you get it now? A3 = 5 + 4 = _? A4 = _ + 4 = _? Np :)